OpenStudy (anonymous):

plz explain it to me, i dont understand :( P(x) is a Polynomial. when P(x) is divided by x-1, the reminder is 4.When P(x) is divided by x+2, the reminder is -5.Find the reminder of P(x) divided by (x-1)(x+2).

OpenStudy (anonymous):

i think you can go about this way, m(x-1) + 4= p(x), n(x+2) -5= p(x).. so what is a(x-1)(x+2) -p(x)=?

OpenStudy (anonymous):

m, n and 'a' being the respective quotients!

OpenStudy (anonymous):

good idea Mashy!

OpenStudy (anonymous):

My textbook says that P(x)=(x-1)(x+2)q(x)+(ax+b) where q(x)is polynomial and a, b are constant how can i get to this step?

OpenStudy (anonymous):

thats just another way of what i said.. !..so you directly go with that